Circular No. 7062

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0111.2–7317 D. Chakrabarty, Massachusetts Institute of Technology; T. Takeshima,

Goddard Space Flight Center, NASA, and Universities Space Research Association; M. Ozaki and B. Paul, Institute of Space and Astronautical Science; and J. Yokogawa, Kyoto University, report: “An observation of the 31-s transient x-ray pulsar XTE J0111.2–7317 with the ASCA satellite on Nov. 18.8–19.6 UT yielded the following improved position for the x-ray source: α = 1^{h}11^{m}14^{s}.5, δ = −73^{o}16′50′′ (equinox J2000.0), with an error radius of 30′′ (90-percent confidence). The x-ray flux in the band 0.7–10 keV was 3.6 × 10^{−10} erg cm^{−2} s^{−1}, corresponding to a luminosity of 2 × 10^{38} erg/s, assuming a distance of 65 kpc. Further observations to identify the optical counterpart are encouraged.”

COMET P/1998 U3 (JÄGER)

Ephemeris extension to IAUC 7040 (elements MPC 32866):

1998 /99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 23 7^{h}00.^{m}26 +41^{o}24.^{′}1 1.558 2.365 135.^{o}3 17.^{o}1 11.2
Dec. 3 6 58.97 +41 01.0 1.455 2.329 144.5 14.2 11.0
13 6 54.21 +40 19.6 1.371 2.295 153.7 10.9 10.8
23 6 46.69 +39 14.2 1.309 2.265 161.8  7.8 10.6
Jan. 2 6 37.84 +37 41.6 1.273 2.238 165.1  6.5 10.5
12 6 29.34 +35 43.9 1.263 2.214 160.4  8.6 10.5
22 6 22.80 +33 28.3 1.278 2.194 151.6 12.3 10.4
Feb. 1 6 19.37 +31 04.7 1.316 2.178 141.8 16.2 10.5
11 6 19.50 +28 42.4 1.375 2.165 132.2 19.7 10.5
21 6 23.16 +26 27.4 1.451 2.157 123.2 22.6 10.6
Mar. 3 6 30.04 +24 22.2 1.540 2.153 114.9 24.7 10.8

Total visual magnitude estimates: Nov. 15.55 UT, 11.6 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 19.88, 11.0 (K. Hornoch, Lelekovice, Czech Republic, 0.13-m reflector); 21.16, 11.3 (G. Comello, Roden, The Netherlands, 0.31-m reflector); 23.44, 11.6 (N. Biver, Oahu, HI, 0.26-m reflector); 25.22, 11.2 (J. Bortle, Stormville, NY, 0.41-m reflector); 27.25, 11.3 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector).

1998 November 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green